How can I use R to solve this question?

Chris owns a laser pointer that is powered by two AAAA batteries. A pair of batteries will power the pointer for an average of five hours use, with a standard deviation of 30 minutes. Chris decides to take advantage of a sale and buys 20 2-packs of AAAA batteries. What is the probability that he will get to use his laser pointer for at least 105 hours before he needs to buy more batteries?
Answer #1

Ans:

sampling distribution of total usage hrs:

mean=20*5=100

standard deviation=sqrt(20)*0.5=2.236

z=(105-100)/2.236

z=2.24

P(z>=2.24)=0.0125

or use pnorm function of R